Suppose in two local counties, the locals have toad collecting contests. In the two largest, the...

  1. Suppose in two local counties, the locals have toad collecting contests. In the two largest, the average number of toads collected by contestants is μ1 = 43 with σ1 = 11 and μ2 = 55 (Obviously a county with a larger toad population) with σ2 = 15. Cindy entered into contest number one and collected 56 toads and her sister Mindy entered into contest number two and collected 73 toads. Which sister did better in the contest?
Answer #1

sister 1
z = (56 - 43)/11 = 1.18

sister 2
z = (73 - 55)/15 = 1.2

Sister 2 i.e. Mindy did better than Cindy